Analyzing Manufacturing Cost Accounts Clapton Company manufactures custom guitars in a wide variety of styles. The fol...

  1. Analyzing Manufacturing Cost Accounts

    Clapton Company manufactures custom guitars in a wide variety of styles. The following incomplete ledger accounts refer to transactions that are summarized for May:

    Materials
    May 1 Balance 105,600 May 31 Requisitions (a)
    31 Purchases 500,000
    Work in Process
    May 1 Balance (b) May 31 Completed jobs (f)
    31 Materials (c)
    31 Direct labor (d)
    31 Factory overhead applied (e)
    Finished Goods
    May 1 Balance 0 May 31 Cost of goods sold (g)
    31 Completed jobs (f)
    Wages Payable
    May 31 Wages incurred 396,000
    Factory Overhead
    May 1 Balance 26,400 May 31 Factory overhead applied (e)
    31 Indirect labor (h)
    31 Indirect materials 15,400
    31 Other overhead 122,500

    In addition, the following information is available:

    1. Materials and direct labor were applied to the following jobs in May:
      Job No. Style Quantity Direct Materials Direct Labor
      101 AF1 330 $ 82,500 $ 59,400
      102 AF3 380 105,400 72,600
      103 AF2 500 132,000 110,000
      104 VY1 400 66,000 39,600
      105 VY2 660 118,800 66,000
      106 AF4 330 66,000 30,800
      Total 2,600 $570,700 $378,400
    2. Factory overhead is applied to each job at a rate of 50% of direct labor cost.
    3. The May 1 Work in Process balance consisted of two jobs, as follows:
      Job No. Style Work in Process,
      May 1
      101 AF1 $26,400
      102 AF3 46,000
      Total $72,400
    4. Customer jobs completed and units sold in May were as follows:
      Job No. Style Completed
      in May
      Units Sold
      in May
      101 AF1 X 264
      102 AF3 X 360
      103 AF2 0
      104 VY1 X 384
      105 VY2 X 530
      106 AF4 0

    Required:

    1. Determine the missing amounts associated with each letter by completing the table below. If required, round your unit cost to two decimal places. If an answer is zero, enter in "0". Enter amounts as positive numbers.

    Job No. Style Quan-
    tity
    May 1
    Work in
    Process
    Direct Materials Direct Labor Factory Overhead Total Cost Unit Cost Units Sold Cost of Goods Sold
    No. 101 AF1 330 $ $82,500 $59,400 $ $ $ $
    No. 102 AF3 380 105,400 72,600
    No. 103 AF2 500 132,000 110,000
    No. 104 VY1 400 66,000 39,600
    No. 105 VY2 660 118,800 66,000
    No. 106 AF4 330 66,000 30,800
    Total 2,600 $ $570,700 $378,400 $ $ $

    a. Materials Requisitions $

    b. Work in Process Beginning Balance $

    c. Direct Materials $

    d. Direct Labor $

    e. Factory overhead applied $

    f. Completed jobs $

    g. Cost of goods sold $

    h. Indirect labor $

    2. Determine the May 31 balances for each of the inventory accounts and factory overhead. Enter all amounts as positive numbers.

    Materials $
    Work in Process $
    Finished Goods $
    Factory Overhead $
Answer #1

Job No

Quantity

(A)

May 1 Work in process

Direct materials

Direct labor

Factory overhead

Total cost

(B)

Unit cost

(B)/(A)

Unit sold

Cost of goods sold

101

330

$26,400

$82,500

$59,400

$29,700

$198,000

$600

264

$158,400

102

380

46,000

105,400

72,600

36,300

260,300

685

360

246,600

103

500

132,000

110,000

55,000

297,000

0

0

0

104

400

66,000

39,600

19,800

125,400

313.5

384

120,384

105

660

118,800

66,000

33,000

217,800

330

530

174,900

106

330

66,000

30,800

15,400

112,200

0

0

0

Total

2,600

$72,400

$570,700

$378,400

$189,200

$1,210,700

$700,284

a. Material requisition = Total Direct material (as calculated above)+ Indirect material

= $570,700 + 15,400 = $586,100

b. Work in process beginning balance = $72,400 (as calculated above)

c. Direct material = $570,700

d. Direct labor = $378,400

e. Factory overhead applied = $189,200

f. Completed jobs = Job no. 101 + Job no. 102 + Job no. 104 + Job no. 105

= $198,000 + 260,300 + 125,400 + 217,800 = $801,500

g. Cost of goods sold = $700,284

h. Indirect labor = Wages incurred - Direct labor

= $396,000 - 378,400 = $17,600

2.

Computation of May 31 balances for each of the following:

Material = Beginning balance + purchases - requisition

= $105,600 + 500,000 - 586,100 = $19,500

Work in process = Job 103 + Job 106

= $297,000 + 112,200 = $409,200

Finished goods = Completed jobs - Cost of goods sold

= $801,500 - 700,284 = $101,216

Factory overhead =( Beginning balance + Indirect labor + indirect materials + other overhead ) - Factory overhead applied

=( $26,400 + 17,600 + 15,400 + 122,500) - 189,200

= $181,900 - 189,200

= $7,300 (over applied factory overhead)